Solve the following radical equations:
(((3^6x+4 -2=0)))
3- is the cubed root
6x+4 is under the radical(square root box)
I'm having trouble solving

Answer by stanbon(75887)

3^6x+4 -2=0)))
3- is the cubed root
----------
(6x+4)^(1/3) -2 = 0
(6x+4)^(1/3)=2
Cube both sides to get:
6x+4=8
6x=8
x=4/3
Cheers,
Stan H.